The motion caused by uniformly accelerated motion is expressed by the following equations: vr vi+ at sesi +ut 1/at2 Here, vis the final velocity) Vis the initial velocityl ais the acceleration Sis the final displacement sHuS the initial displacement anduis time Projectile Exercises In each of the following ignore the effects of air resistance on projectile motion 1) In soccer, a dropped-ball is a procedure used to restart play if play is stopped and neither team was in clear possession of the ball. The rules state that the referee drops the ball from waist height between opposing players, who can play the ball once it has struck the ground. Howard Webb was an English FIFA referee who refereed the 2010 World Cup Final. With a height of 1.88 m Howard Webbs waist height would be approximately 1 m. If he were to restart play with a dropped-ball, how long would it take for the ball to strike the ground and how fast would it be going when it struck the ground?
Answer #1

1) time taken to reach to the ground  

S = ut+0.5gt^2

1.88 = 0+0.5*9.8*t^2

t = 0.62 s

speed when it hits the ground

v^2 - u^2 = 2gS

v^2 - 0 = 2*9.8*1.88

v = 6.07 m/s